# Don’t Let It Hallucinate: Premise Verification via Retrieval-Augmented Logical Reasoning

### v1

@@ A.2 Case Study
Logical Form:
is a recipient of(’The Dark Knight’, 16th Screen Actors Guild Awards)
Retrieved Graph Triple:
[’The Dark Knight’, ’award received’, ’81st Academy Awards’]
@@
Corrected Answer: No, "The Dark Knight" was not the recipient of the 16th Screen Actors Guild Awards. The awards ceremony took place in 2010, and "The Hurt Locker" won the award for Outstanding Performance by a Cast in a Motion Picture.
A.3 Prompt Details
The following prompt is used to combine the information retrieved from the knowledge graph $G$ (context) and the query logical form $\mathcal{L}(q)$ (query) to form the input to the LLMs discussed in §4.2.1.

### v2

@@ A.4.6 Case Study
Logical Form: is a recipient of(’The Dark Knight’, 16th Screen Actors Guild Awards)
Retrieved Graph Triple: [’The Dark Knight’, ’award received’, ’81st Academy Awards’]
@@
Corrected Answer:
No, "The Dark Knight" was not the recipient of the 16th Screen Actors Guild Awards. That year’s SAG Award for Outstanding Performance by a Cast in a Motion Picture went to Inglourious Basterds, not The Dark Knight.
A.5 Evaluation under Longer Queries
Longer questions often introduce additional contextual cues, compositional structures, and distracting details, which can obscure the core false assumption and make premise verification more challenging. Evaluating under such settings is important for understanding whether a model can robustly identify and reason about incorrect premises rather than relying on surface-level patterns.
As an initial step, we select the first 105 questions from the KG-FPQ dataset and use an LLM (gpt-4o-mini) to rewrite them into longer, more natural user queries while preserving the original (true or false) premises. We then evaluate our method on both false-premise detection and hallucination mitigation under this longer-context setting. The prompt used for query rewriting, along with the rewritten queries, is released in accompanying repository [^4] https://github.com/yqin43/premise-verification. The False Premise Detection result and Hallucination Mitigation result are shown in Tab. 19 and Tab. 20. Overall, the results on this longer query subset show that the proposed method consistently improves performances on both false-premise detection and hallucination mitigation tasks, suggesting that its effectiveness extends beyond short or minimally phrased queries.
Metric Direct Claim with RAG G-retriever GraphRAG/ToG
Original Query for Both Stages
TPR 2.22 85.56 68.89 92.22
TNR 50.00 93.33 73.33 53.33
FPR 50.00 6.67 26.67 46.67
FNR 97.78 14.44 31.11 7.78
F1 4.04 91.67 79.49 92.22
Accuracy 8.65 86.67 69.52 86.67
Logical Form for Retrieval and Original Query for Detection
TPR 2.22 94.44 93.33 92.22
TNR 50.00 73.33 53.33 53.33
FPR 50.00 26.67 46.67 46.67
FNR 97.78 5.56 6.67 7.78
F1 4.04 94.97 92.82 92.22
Accuracy 8.65 91.43 87.62 86.67
Logical Form for Both Stages
TPR 2.22 94.44 97.78 92.22
TNR 50.00 73.33 46.67 53.33
FPR 50.00 26.67 53.33 46.67
FNR 97.78 5.56 2.22 7.78
F1 4.04 94.97 94.62 92.22
Accuracy 8.65 91.43 90.48 86.67
Table 19: False-premise detection performance under longer queries.
Setting Accuracy
Direct Ask 0.933
Ours 0.952
Table 20: Hallucination mitigation accuracy under longer queries using GPT-4o-mini.
A.6 Logical Form Correctness Validation
Our current human evaluation was conducted by two Ph.D. students with relevant NLP/LLM research experience. All annotators reached full agreement, resulting in an inter-annotator agreement of 1.0.
A.7 Additional Experiment Setup
All models are implemented and run on a multi-NVIDIA RTX 6000 Ada workstation. For Logical Form Extraction and Retrieval, we set parameters temperature =0=0 and top_p = 1.